hallm:

Solve for x. 38(2x+16)−2=13 Enter your answer in the box

Razor:

is 38 supposed to be a fraction or naw?

hallm:

yes is it

563blackghost:

So it is... \(\large\bf{\frac{3}{8} (2x+16)-2=13}\) ?

Razor:

^^

hallm:

yes

563blackghost:

Okay. First you need to distribute the \(\bf\frac{3}{8}\) into the parenthesis. \(\Large\bf{\frac{3}{8} (2x+16) ~~\rightarrow ~~~~(\frac{3}{8} \times 2x)+(\frac{3}{8} \times 16)}\) What does that simplify to?

hallm:

thats what i dont know i need help figuring it out

563blackghost:

I know im telling you how. You distribute the fraction into the parenthesis. What does \(\large\bf{\frac{3}{8} \times 2x}\) equal?

hallm:

4/16

563blackghost:

Not quite. \(\large\bf{\frac{3}{8} \times 2x ~~~ \rightarrow \color{red}{\frac{3 \times 2}{8}x}}\) What does that simplify to?

hallm:

6/8

563blackghost:

correct. now this can be simplified further. \(\large\bf{\frac{6}{8} \rightarrow \frac{3}{4}}\) by division of 2. Now we distribute the fraction to 16. \(\Large\bf{\frac{3}{8} \times 16 \rightarrow \frac{3 \times 16}{8}}\)

hallm:

48/8

563blackghost:

Okie now simplify that

hallm:

6

563blackghost:

Correct. So our distributed parenthesis is simplified to: \(\Large\bf{\frac{3}{8}(2x+16) \rightarrow \color{red}{\frac{3}{4}x+6}}\) So we have: \(\Large\bf{\frac{3}{4}x+6-2=13}\) Got it?

hallm:

yayyyyy thank you so so so much

563blackghost:

not done yet

563blackghost:

still have to simplify further

hallm:

ok what do i do now

563blackghost:

Now you need to subtract 6 and 2. \(\Large\bf{\frac{3}{4}x+\color{red}{6-2} ~~\rightarrow \frac{3}{4}+4=13}\)

563blackghost:

3/4x*

hallm:

4

563blackghost:

Now you need to get the variable to one side. You see that `4` is being added to `3/4x` this means you need to `SUBTRACT` `4` from both sides to get it to one side. \(\Large\bf{\frac{3}{4}x + 4 \color{red}{-4} = 13 \color {red}{-4}}\) What does that simplify to?

hallm:

9

563blackghost:

Correct. So we now have: \(\Large\bf{\frac{3}{4}x=9}\) Now you need to apply the fraction to the other side. So you divide the fraction to get x to itself. \(\huge\bf{\frac{\frac{3}{4}}{\color{red}{\frac{3}{4}}}x=\frac{9}{\color{red}{\frac{3}{4}}}}\) What does x equal?

hallm:

ummmm

563blackghost:

To simplify by fraction you follow by: \(\large\bf{\frac{a}{\frac{b}{c}}=\frac{a \times c}{b}}\) So: \(\huge\bf{x=\frac{9 \times 4}{3}}\)

hallm:

32/3?

hallm:

12

hallm:

x=12

563blackghost:

Correct! So `x=12` cx

563blackghost:

nice job!

hallm:

will you be my tutor

563blackghost:

cx uh well if you ever need help just dm me >.< i actually have to go right now ;-;

hallm:

ok thank you